Write the expression in the form a + bi, where a and b are real numbers.
__9___
6 + 2i

Answer by shweta(56)

__9___
6 + 2i
rationalise it
__9___ * _(6-2i)_
6 + 2i * (6-2i)
__9*(6-2i)__
(6 + 2i)*(6-2i)
Apply formula:(a+b)*(a-b)=a^2 -b^2
__54-18i__
36-4i^2
Now i^2= -1
__54-18i__
36-4*(-1)
__54-18i__
36+4
__54-18i__
40
54/40 -18i/40